OpenStudy (anonymous):

The sum of two numbers is 62 and their difference is 8. What are the numbers? Please put in equation form

OpenStudy (agl202):

if x + y = 62 , and |x - y| = 8, there are a couple of options. If x is the larger number, we could sub in "y +8" for x, and the equation would look like this: y + ( y + 8 ) = 62 >> 2y + 8 = 62 >> 2y = 54 >> y = 27, and x = 27 + 8 or 35. If we switch it around, so that y is the larger number, you get y = 35, x = 27. so there are two different possibilities 1) y = 27, x = 35 2) x = 27, y = 35 Hope this helps! : )